A rectangular field is 300 feet wide and 400 feet long. Random sampling indicates that on the averagre, three ants can be fond per square inch throughout the field. Find the number of ants in the field?
Answer by checkley71(8403)

300*400=120,000 sq ft.
120,000*144=17,280,000 sq inches.
17,280,000*3=51,840,000 ants in the field.